Square root of 12x^3
Answer by stanbon(48535)

Square root of 12x^3
Break 12x^3 into two factors making the 1st factor
the largest possible perfect square: for example.
12x^3 = (4x^2)(3x)
Then sqrt[12x^3]=sqrt[(4x^2)(3x)]
Rewrite as [sqrt(4x^2)][sqrt(3x)]
Since the 4x^2 is a perfect square you can take the square root and get:
2x[sqrt(3x)]
That's as far as you can go with this problem.
So sqrt(12x^3)= 2x[sqrt(3x)]
Cheers,
stan H.